OpenStudy (anonymous):

A recreation room has the following dimensions. Carpeting will cost $9.00 a square yard. How many square yards of carpet will be needed and how much will it cost? Include correct units at each stage within your solution.

OpenStudy (anonymous):

First you need to find the area of the different sections of the room. There is a large rectangle of 6 x 15 yards => 90 sq. yards The triangle has the same height as the large rectangle and the base is the same as the square below it. area for triangle = 1/2 * b* h => 1/2 * 6 * 4 = 12 sq. yards The square is 4*4 = 16 sq. yards. Add the areas : 90 + 12 + 16 = 118 sq. yards. Multiply the number of sq. yards by the price per sq. yard : 118 * 9.00 = 1062 So you need 118 sq. yards of carpet, which will cost $1,062

OpenStudy (anonymous):